# The lengths of conjugators in the model filiform groups

### v1

@@ 4. Centralisers in $\Gamma_{d}$
Picking up a theme from the last argument of the previous section, we note that $g\in\Gamma_{d}$ will conjugate $\gamma$ to an element of the form $\gamma a_{d}^{*}$ if and only if the image of $g$ in $\Gamma_{d-1}$ centralises the image of $\gamma$. This highlights our need to understand centralisers in the groups $\Gamma_{d}$.
@@
    $C(\gamma)=\langle\gamma_{0},a_{d}\rangle\cong\mathbb{Z}^{2}$ if $\gamma_{0}\not\in{\mathbb{A}}_{d}$,
@@ 5. The lengths of roots in $\Gamma_{d}$
Before proving Proposition 5.1, we consider the case $d=2$ in order to elucidate the geometry of what is happening.
@@
$\sigma(n)=1+\dots+n=n(n-1)/2={\rm{Area}}_{\langle a_{1},t\rangle}(u^{n}v^{-1})$
@@
$(a_{1}ta_{2}^{(1-n)/2})^{n}=t^{n}a_{1}^{n}.$
Thus $h_{n}:=a_{1}ta_{2}^{(1-n)/2}$ is the unique $n$-th root of $g:=t^{n}a_{1}^{n}$ in $\Gamma_{2}$. Projecting to $\Gamma_{1}=\mathbb{Z}^{2}$, it is easy to see that $d(1,t^{n}a_{1}^{n})=2n$ in $\Gamma_{2}$. On the other hand, $d(1,h_{n})$ is approximately $\sqrt{n}$, even though $h_{n}^{n}=g$.
@@
where $|r|\leq{\rm{Area}}_{\Gamma_{1}}(v^{-p}{\overline{w}})$, which is less than $n^{2}$, by the quadratic isoperimetric inequality in $\mathbb{Z}^{2}$. (In more detail, $v^{-p}{\overline{w}}$ is a non-empty word $\tau$ in $a_{1}$ and $t$ of length at most $2n$ that equals the identity in $\Gamma_{1}$ and any such $\tau$ has a subword which freely reduces to $a_{1}^{\pm 1}t^{l}a_{1}^{\mp 1}$ or $t^{\pm 1}a_{1}^{l}t^{\mp 1}$ for some $|l|\leq n/2$; this subword can be replace by $t^{l}$ or $a_{1}^{l}$, respectively, at a cost of applying at most $n/2$ commutator relations, from which it follows that ${\rm{Area}}(\tau)<n^{2}$, by an induction on length.) But $v=ha_{2}^{j}$ in $\Gamma_{2}$ for some $j\in\mathbb{Z}$, so
@@
We saw earlier that $|v|\leq n/p$ and $|r|{<}n^{2}$, and from Proposition 3.2 we know $d(1,a_{2}^{-j})\leq C_{2}\,\sqrt{j}=C_{2}\sqrt{r/p}$, which is less than $C_{2}\,n/\sqrt{p}$. Thus, writing $v\in\langle a_{1},t\rangle$ in normal form $v=t^{m_{0}}a_{1}^{m_{1}}$ with $|m_{0}|+|m_{1}|\leq n/p$, we have $h=va_{2}^{-j}=t^{m_{0}}a_{1}^{m_{1}}a_{2}^{{-j}}$ in the form required by Proposition 5.1, since $|j|\leq n^{2}/p$. Moreover,
@@
We proceed by induction on $d$. The case $d=1$ is trivial and the case $d=2$ was covered in the preceding example. Assume true up to $d-1$. We have $\gamma,h\in\Gamma_{d}$ with $h^{p}=\gamma$ and $d(1,\gamma)\leq n$. Let ${\overline{h}}$ be the image of $h$ in $\Gamma_{d-1}$. Then, by induction, ${\overline{h}}=t^{m_{0}}a_{1}^{m_{1}}\dots a_{d-1}^{m_{d-1}}$, in $\Gamma_{d-1}$, with the $|m_{i}|$ bounded as stated in the proposition. Let
@@
then we will have $|mp|\leq(\alpha_{d}+1)n$ and defining $k_{d}:=\max\left\{\alpha_{d}+1,k_{d-1}\right\}$ will complete the induction.
@@
6. On the geometry of numbers
We need an elementary observation concerning Bézout’s Lemma. Consider positive integers $A\neq B$ where $A$ does not divide $B$. Let $e=\gcd(A,B)$ and observe that the points $(\mu,\lambda)\in\mathbb{R}^{2}$ with $\lambda A-\mu B=e$ form a line $L$ parallel to the vector $(A,B)$ that includes the points $\mathbf{p}=(0,e/A)$ and $\mathbf{q}=(A,(e/A)+B)$. Let $S=[\mathbf{p},\mathbf{q})$ be the line-segment between these two points, but excluding $\mathbf{q}$. Bézout’s Lemma tells us that there is a point $\mathbf{x}=(\lambda_{0},\mu_{0})$ on $L$ with integer coefficients; translating $\mathbf{x}$ by a suitable integer multiple of $(A,B)$, we may assume that it lies on $S$. So $0\leq\mu_{0}<A$ and $e/A\leq\lambda_{0}<(e/A)+B$. But $\mu_{0}\neq 0$, because $\lambda_{0}A$ cannot equal $e$ when $\lambda_{0}$ is an integer. So $0<\mu_{0}<A$. And, because $1\leq e<A$, the inequalities constraining $\lambda_{0}$ imply that $0<\lambda_{0}<1+B$. As $\lambda_{0}$ and $B$ are integers, it follows that $\lambda_{0}\leq B$.
@@ 6. On the geometry of numbers
For positive integers $A\neq B$, if $e=\gcd(A,B)<A$, then there exist positive integers $\mu<A$ and $\lambda\leq B$ such that $\lambda A-\mu B=e$.
@@
and these are equal if and only if $s\alpha+\eta=r\omega+\xi$. This condition can be expressed as $s\alpha-r\omega=\xi-\eta$, and if we regard this as an equation with variables $s$ and $\omega$, Lemma 6.1 implies that this equation has an integer solution with $|s|$ and $|\omega|$ both at most $\max\left\{|\alpha|,|r|,|\xi-\eta|\right\}$. But, by Lemma 3.1, $|\alpha|$ and $|r|$ are at most $n$, and $|\xi|$ and $|\eta|$ are at most $n^{2}$. The existence of a suitable constant $B_{2}$ follows.
@@
Theorem 2.2 provides an integer $A_{d-1}$ and such that
${\rm{Area}}_{\Gamma_{d-1}}(w^{-1}{\overline{u}}w{\overline{v}}^{-1})\leq A_{d-1}|w^{-1}{\overline{u}}w{\overline{v}}^{-1}|^{d}\leq A_{d-1}(2n+{\color[rgb]{0.80,0.68,0.00}2}B_{n-1}n^{d-1})^{d}\leq A_{d-1}^{\prime}n^{d(d-1)},$
where for convenience we define $A_{d-1}^{\prime}:=2^{d}A_{d-1}(1+B_{n-1})^{d}$. It follows from Lemma 2.1 that $w^{-1}{u}w={v}a_{d}^{\ell}$ in $\Gamma_{d}$ where $|\ell|\leq A_{d-1}^{\prime}n^{d(d-1)}$.
@@
Claim 1: If $g\in\Gamma_{d-1}\smallsetminus{\mathbb{A}}_{d-1}$ with $d(1,g)\leq n$ and $\ell\in{\rm{im}\,}\zeta_{g}$ with $|\ell|\leq A\,n^{d(d-1)}$, then there exists $y\in C_{\Gamma_{d-1}}(g)$ with $d(1,y)<AE_{d}\,n$ and $\zeta_{g}(y)=\ell$.
@@
$\ell=Mpq+\rho qe$
for integers $\rho$ and $M$ with $0\leq\rho<p$ and $|M|\leq A\,n^{d(d-1)}$. Thus, using (4) and (5), we have
$\ell=M\,\zeta_{g}(a_{d-1}^{-1})+\rho\,\zeta_{g}(g_{0}^{\lambda}a_{d-1}^{\mu})=\zeta_{g}(a_{d-1}^{-M}(g_{0}^{\lambda}a_{d-1}^{\mu})^{\rho}).$
@@
$d(1,y)=d(1,a_{d-1}^{-M}(g_{0}^{\lambda}a_{d-1}^{\mu})^{\rho})$ $\leq d(1,a_{d-1}^{M})+\rho\,d(1,g_{0}^{\lambda}a_{d-1}^{\mu})$

### v2

@@ 4. Centralisers in $\Gamma_{d}$
Picking up a theme from the last argument of the previous section, we note that $g\in\Gamma_{d}$ will conjugate $\gamma$ to an element of the form $\gamma a_{d}^{*}$ if and only if the image of $g$ in $\Gamma_{d-1}$ centralises the image of $\gamma$. This highlights our need to understand centralisers $C(\gamma)$ in the groups $\Gamma_{d}$.
@@
    $C(\gamma)=\langle\gamma_{0},a_{d}\rangle\cong\mathbb{Z}^{2}$ if $\gamma\not\in{\mathbb{A}}_{d}$,
@@ 5. The lengths of roots in $\Gamma_{d}$
Before proving Proposition 5.1, we consider the case $d=2$, which admits a more direct argument.
@@
$\sigma(n)=1+\dots+n=n(n+1)/2={\rm{Area}}_{\langle a_{1},t\rangle}(u^{n}v^{-1})$
@@
$(a_{1}ta_{2}^{-(n+1)/2})^{n}=t^{n}a_{1}^{n}.$
Thus $h_{n}:=a_{1}ta_{2}^{-(n+1)/2}$ is the unique $n$-th root of $g:=t^{n}a_{1}^{n}$ in $\Gamma_{2}$. Projecting to $\Gamma_{1}=\mathbb{Z}^{2}$, it is easy to see that $d(1,t^{n}a_{1}^{n})=2n$ in $\Gamma_{2}$. On the other hand, $d(1,h_{n})$ is approximately $\sqrt{n}$, even though $h_{n}^{n}=g$.
@@
where $|r|\leq{\rm{Area}}_{\Gamma_{1}}(v^{-p}{\overline{w}})$, which is less than $n^{2}$, by the quadratic isoperimetric inequality in $\mathbb{Z}^{2}$. (In more detail, $v^{-p}{\overline{w}}$ is a non-empty word $\tau$ in $a_{1}$ and $t$ of length at most $2n$ that equals the identity in $\Gamma_{1}$ and any such $\tau$ has a subword which freely reduces to $a_{1}^{\pm 1}t^{l}a_{1}^{\mp 1}$ or $t^{\pm 1}a_{1}^{l}t^{\mp 1}$ for some $|l|\leq n/2$; this subword can be replaced by $t^{l}$ or $a_{1}^{l}$, respectively, at a cost of applying at most $n/2$ commutator relations, from which it follows that ${\rm{Area}}(\tau)<n^{2}$, by an induction on length.) But $v=ha_{2}^{j}$ in $\Gamma_{2}$ for some $j\in\mathbb{Z}$, so
@@
We saw earlier that $|v|\leq n/p$ and $|r|{<}n^{2}$, and from Proposition 3.3 we know $d(1,a_{2}^{-j})\leq C_{2}\,\sqrt{|j|}=C_{2}\sqrt{|r/p|}$, which is less than $C_{2}\,n/\sqrt{p}$. Thus, taking $v\in\langle a_{1},t\rangle$ to be $v=t^{m_{0}}a_{1}^{m_{1}}$ with $|m_{0}|+|m_{1}|\leq n/p$, we have $h=va_{2}^{-j}=t^{m_{0}}a_{1}^{m_{1}}a_{2}^{{-j}}$ in the form required by Proposition 5.1, since $|j|\leq n^{2}/p$. Moreover,
@@
We proceed by induction on $d$. The case $d=1$ is trivial and the case $d=2$ was covered in the preceding example. Assume true up to d1d-1. We have $\gamma,h\in\Gamma_{d}$ with $h^{p}=\gamma$ and $d(1,\gamma)\leq n$. Let $\overline{h}$} be the image of $h$ in $\Gamma_{d-1}$. Then, by induction, ${\overline{h}}=t^{m_{0}}a_{1}^{m_{1}}\dots a_{d-1}^{m_{d-1}}$, in $\Gamma_{d-1}$, with the $|m_{i}|$ bounded as stated in the proposition. Let
@@
then we will have $|mp|\leq(\alpha_{d}+1)n^{d}$ and defining $k_{d}:=\max\left\{\alpha_{d}+1,k_{d-1}\right\}$ will complete the induction.
@@
6. Quantifying Bézout’s Lemma
We will need an elementary observation concerning Bézout’s Lemma.
@@ 6. Quantifying Bézout’s Lemma
For positive integers $A$ and $B$, if $e=\gcd(A,B)$ then there exist a non-negative integer $\mu<A$ and a positive integer $\lambda\leq B$ such that $\lambda A-\mu B=e$. If $A>1$, one can assume $\mu$ is positive.
Proof.
By Bézout’s Lemma, there exist non-zero integers $a$ and $b$ such that $aA+bB=e$. If $A=1$, take $\lambda=1$ and $\mu=0$. If $A>1$ and $B=kA$, take $\lambda=k+1$ and $\mu=1$. In all other cases, $A$ does not divide $e$, so it does not divide $b$, and b=kArb=kA-r with $0<r<A$. Then (kB+a)ArB=e(kB+a)A-rB=e and
$(kB+a)=(r/A)B+(e/A)$
is a positive integer less than or equal to $B$, because $r/A$ and $e/A$ are both positive numbers less than $1$. Let $\lambda=(kB+a)$ and let $\mu=r$. ∎
The simple estimate in the following corollary can be sharpened (cf. [BR25a, Corollary 3.3]) but it is sufficient for our needs in this article.
Corollary 6.2.
Let $A,B$ and $m$ be integers and let $N=\max\{|A|,\,|B|\}$. If there is an integer solution $(x,y)$ to the equation $Ax+By=m$, then there is an integer solution $(x_{0},y_{0})$ with $\max\{|x_{0}|,\ |y_{0}|\}\leq|m|+N^{2}$.
Proof.
If $N=0$ the assertion is trivial, so suppose $|A|\geq 1$. Write $m=kA+r$ with $|k|\leq|m|$ and $0\leq r<|A|$. As $m$ lies in the subgroup of $\mathbb{Z}$ generated by $A$ and $B$, so does $r$. This subgroup is generated by $e=\gcd(A,B)$, so $r=r_{0}e$ for some $r_{0}\leq r<|A|$. The lemma allows us to write $e=\lambda_{0}A+\mu_{0}B$ with $\max\{|\lambda_{0}|,\,|\mu_{0}|\}\leq N$. Then
$m=(k+\lambda_{0}r_{0})A+\mu_{0}B,$
and $|k+\lambda_{0}r_{0}|\leq|k|+r|\lambda_{0}|\leq|m|+N^{2}$. ∎
@@
and these are equal if and only if $s\alpha+\eta=r\omega+\xi$. This condition can be expressed as $s\alpha-r\omega=\xi-\eta$, and if we regard this as an equation with variables $s$ and $\omega$, Corollary 6.2 implies that this equation has an integer solution with $|s|$ and $|\omega|$ both at most $|\xi-\eta|+\max\left\{|\alpha|^{2},|r|^{2}\right\}$. Lemma 3.2 tells us that $|\alpha|$ and $|r|$ are at most $n$ and $|\xi|$ and $|\eta|$ are at most $n^{2}$. The existence of a suitable constant $B_{2}$ follows.
@@
Theorem 2.2 provides an integer $A_{d-1}$ such that
${\rm{Area}}_{\Gamma_{d-1}}(w^{-1}{\overline{u}}w{\overline{v}}^{-1})\leq A_{d-1}|w^{-1}{\overline{u}}w{\overline{v}}^{-1}|^{d}\leq A_{d-1}(2n+{2}B_{d-1}n^{d-1})^{d}\leq A_{d-1}^{\prime}n^{d(d-1)},$
where for convenience we define $A_{d-1}^{\prime}:=2^{2d}A_{d-1}B_{d-1}^{d}$. It follows from Lemma 2.1 that $w^{-1}{u}w={v}a_{d}^{\ell}$ in $\Gamma_{d}$ where $|\ell|\leq A_{d-1}^{\prime}n^{d(d-1)}$.
@@
Claim 1: If $g\in\Gamma_{d-1}\smallsetminus{\mathbb{A}}_{d-1}$ with $d(1,g)\leq n$ and $\ell\in{\rm{im}\,}\zeta_{g}$ with $|\ell|\leq A\,n^{d(d-1)}$, then there exists $y\in C_{\Gamma_{d-1}}(g)$ with $d(1,y)<AE_{d}\,n^{d}$ and $\zeta_{g}(y)=\ell$.
@@
$\ell=Mpq+\theta qe$
for integers $\theta$ and $M$ with $0\leq\theta<p$ and $|M|\leq A\,n^{d(d-1)}$. Thus, using (4) and (5), we have
$\ell=M\,\zeta_{g}(a_{d-1}^{-1})+\theta\,\zeta_{g}(g_{0}^{\lambda}a_{d-1}^{\mu})=\zeta_{g}(a_{d-1}^{-M}(g_{0}^{\lambda}a_{d-1}^{\mu})^{\theta}).$
@@
$d(1,y)=d(1,a_{d-1}^{-M}(g_{0}^{\lambda}a_{d-1}^{\mu})^{\theta})$ $\leq d(1,a_{d-1}^{M})+\theta\,d(1,g_{0}^{\lambda}a_{d-1}^{\mu})$